OpenStudy (anonymous):

Write the expression as the sine, cosine, or tangent of an angle. cos 96° cos 15° + sin 96° sin 15°

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Hmm so this can be figured out using the `Angle Sum Formula for Cosine`\[\large \cos(\alpha-\beta)=\cos \alpha \cos \beta+\sin \alpha \sin \beta\]